OpenStudy (anonymous):

I need to know if I got this right or not. Write the equation of the line that passes through (–2, 6) and (2, 14) in slope-intercept form. y = 2x – 2 << my answer but I'm not sure y = 2x + 10 y = 0.5x + 7 y = 0.5x – 4

OpenStudy (sohailiftikhar):

yes your answer is correct :)

OpenStudy (anonymous):

thank you

OpenStudy (sohailiftikhar):

np :)

OpenStudy (mathstudent55):

Start with the slope-intercept form of the equation of a line. Then plug in the given points. This is how you do it. \(y - y_1 = \dfrac{y_2 - y_1}{x_2 - x_1}(x - x_1) \) \(y - 6 = \dfrac{14 - 6}{2 - (-2)}(x - (-2)) \) \(y - 6 = \dfrac{8}{4}(x + 2) \) \(y - 6 = 2(x + 2) \) \(y - 6 = 2x + 4) \) \(y = 2x + 10 \) The answer is choice B. To check, you can confirm that both points work in the equation of choice B. Choice A is not correct. In fact, neither point works in the equation of choice A.

OpenStudy (anonymous):

Oops yes B is correct

OpenStudy (sohailiftikhar):

oh yes B is correct sorry .